OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (anonymous):

@billj5 I can't figure out how to set it up

OpenStudy (jdoe0001):

hint: dot product \(\bf <a,b,c>\cdot <d,e,f>\implies (a\cdot d)+(b\cdot e)+(c\cdot f)\)

OpenStudy (anonymous):

v=6i+7j-3k w=-7i+5j+2k \[v.w=(6i+7j-3k).(-7i+5j+2k)=6(-7)+7(5)-3(2)=?\]